# Adapting LLaMA Decoder to Vision Transformer

### v1

@@ 4.2 Comparison with Recent Architectures on ImageNet-1K Classification
Results. Table 6 shows a comparison between iLLaMA and other strong visual baselines, including ConvNets (ConvNeXt liu2022convnet , ConvNeXt-V2 woo2023convnext ), vision Transformers (ViT dosovitskiy2020image , Swin Transformer liu2021swin ), MLPs (PoolFormer yu2022metaformer , VanillaNet chen2023vanillanet ), and recent language model inspired models (AIM el2024scalable , VisionLLaMA chu2024visionllama ). We present three observations: 1) The performance-parameter trade-off of iLLaMA surpasses other LM-inspired models such as AIM and VisionLLaMA, presumably due to its use of casual attention and soft mask training techniques. 2) iLLaMA exhibits a superior accuracy-throughput trade-off compared to strong hierarchical baselines such as ConvNeXt-V2-N/T/B and Swin-S/B. We attribute this to iLLaMA’s isotropic design (each intermediate block has the same feature resolution), which benefits from a straightforward and efficient architecture, enhancing inference speed. 3) Scalability of model capacity and dataset size: After comprehensive pre-training on the expanded ImageNet-21K dataset, the iLLaMA-B model achieves more than $85.0\%$ accuracy on ImageNet-1K with under 100M parameters, significantly outperforming ViT-B’s $84.0\%$. Upon scaling up to the larger iLLaMA-L, accuracy reaches $86.0\%$, exceeding that of ViT-L pre-trained on ImageNet-21K and the AIM-7B pre-trained on the DFN-2B+ dataset. To our knowledge, this showcases SOTA performance for LLaMA-type architectures.
@@ 4.4 Compatibility with Quantization
Since a practical goal for neural networks is deployment on low-bit hardware chips, we further examine iLLaMA’s compatibility with quantization. We basically follow Q-ViT li2022q to apply quantization-aware training (QAT) to iLLaMA, with weights and activations of all blocks’ FFN and casual self-attention layers to 8 bits. Quantization recipes and results are shown in Appendix B.4 and Table 8, respectively. Different sizes of low-bit iLLaMA maintain accuracy well, and 8-bit iLLaMA-T is even compete favorably with the full-precision DeiT-T touvron2021training ($72.4\%$ v.s. $72.2\%$).
@@ 4.5 Transferability on Downstream Tasks
ADE20K semantic segmentation. We fine-tune our ImageNet-1K pre-trained iLLaMA and ViT models on ADE20K zhou2019semantic dataset using UperNet xiao2018unified to perform semantic segmentation task. For both iLLaMA and ViT, we set the learning rate as 6e-5 and weight decay as 0.01. Table 10 presents the results. iLLaMA’s performance is marginally lower than ViT’s, which we attribute to the potential impact of the masking mechanism in iLLaMA’s casual attention on high-resolution dense prediction tasks. This suggests there is still space for optimization, a subject for future investigation.
@@ 5 Conclusions
In the general trend of using encoder to build vision architectures, the suitability for decoder-only architecture still exists a research gap. We systematically studies whether decoder, an architecture that has shown amazing potential in LLMs, can also take root in learning visual representation through visual supervised training. The key component – casual self-attention we used – is not novel and is inherited from existing LLM architectures, but we propose pivotal techniques, i.e., PS [cls] and soft mask, to effectively adapt them to visual tasks. We hope that this work will inspire more exploration towards a generic pure decoder architecture that fully unifies vision and text.
@@ Appendix A Network Configuration
We provide a block-level comparison between iLLaMA and ViT model in Figure 5. VisionLLaMA uses SwiGLU, and AS2D RoPE to build LLaMA-style architecture. Differently, we further uses RMSNorm, modified casual self-attention and 1D RoPE from LLaMA to replace layer normalization, bi-directional self-attention, and proposes two pivotal strategies, i.e., PS [cls] and soft mask to help the optimization of our iLLaMA. We also keep the learnable positional embedding as ViT does.
@@ B.2 ImageNet (Pre-)training Recipe
We pre-trained the iLLaMA-B/L models on ImageNet-21K for 90 epochs, adhering to the practices in liu2022convnet . We set the cutoff epochs to 30, indicating that the iLLaMA models’ self-attention fully transitions to casual self-attention after 30 epochs.
@@ B.4 Quantization-aware Training Recipe
We provide our quantization-aware training recipe for iLLaMA in Table 15. Basically we follow the Q-ViT method proposed in li2022q , with only weights and activations in each basic block’s casual self-attention and FFN module are quantized to 8 bit width.
@@ Appendix C Designing iLLaMA: detailed results
Table 17: ImageNet-1K classification accuracy via gradually replacing components in ViT-T/16 and ViT-B/16 with counterparts in LLaMA, better or worse than the ViT baseline results with our basic training recipe. Components from or modified from LLaMA are highlighted. P.E.: positional embedding, Bd.: bi-directional self-attention, Cs.: casual self-attention.
@@
Appendix D Rank Analysis of Casual Self-attention

### v2

@@ 4.2 Comparison with Recent Architectures on ImageNet-1K Classification
Results. Table 6 shows a comparison between iLLaMA and other strong visual baselines, including ConvNets (ConvNeXt liu2022convnet , ConvNeXt-V2 woo2023convnext ), vision Transformers (ViT dosovitskiy2020image , Swin Transformer liu2021swin ), MLPs (PoolFormer yu2022metaformer , VanillaNet chen2023vanillanet ), and recent language model inspired models (AIM el2024scalable , VisionLLaMA chu2024visionllama ). We present three observations: 1) The performance-parameter trade-off of iLLaMA surpasses other LM-inspired models such as AIM and VisionLLaMA, presumably due to its use of causal attention and soft mask training techniques. 2) iLLaMA exhibits a superior accuracy-throughput trade-off compared to strong hierarchical baselines such as ConvNeXt-V2-N/T/B and Swin-S/B. We attribute this to iLLaMA’s isotropic design (each intermediate block has the same feature resolution), which benefits from a straightforward and efficient architecture, enhancing inference speed. 3) Scalability of model capacity and dataset size: After comprehensive pre-training on the expanded ImageNet-21K dataset, the iLLaMA-B model achieves more than $85.0\%$ accuracy on ImageNet-1K with under 100M parameters, significantly outperforming ViT-B’s $84.0\%$. Upon scaling up to the larger iLLaMA-L, accuracy reaches $86.0\%$, exceeding that of ViT-L pre-trained on ImageNet-21K and the AIM-7B pre-trained on the DFN-2B+ dataset. To our knowledge, this showcases SOTA performance for LLaMA-type architectures.
@@ 4.4 Compatibility with Quantization
Since a practical goal for neural networks is deployment on low-bit hardware chips, we further examine iLLaMA’s compatibility with quantization. We basically follow Q-ViT li2022q to apply quantization-aware training (QAT) to iLLaMA, with weights and activations of all blocks’ FFN and causal self-attention layers to 8 bits. Quantization recipes and results are shown in Appendix B.4 and Table 8, respectively. Different sizes of low-bit iLLaMA maintain accuracy well, and 8-bit iLLaMA-T is even compete favorably with the full-precision DeiT-T touvron2021training ($72.4\%$ v.s. $72.2\%$).
@@ 4.5 Transferability on Downstream Tasks
ADE20K semantic segmentation. We fine-tune our ImageNet-1K pre-trained iLLaMA and ViT models on ADE20K zhou2019semantic dataset using UperNet xiao2018unified to perform semantic segmentation task. For both iLLaMA and ViT, we set the learning rate as 6e-5 and weight decay as 0.01. Table 10 presents the results. iLLaMA’s performance is marginally lower than ViT’s, which we attribute to the potential impact of the masking mechanism in iLLaMA’s causal attention on high-resolution dense prediction tasks. This suggests there is still space for optimization, a subject for future investigation.
@@ 5 Conclusions
In the general trend of using encoder to build vision architectures, the suitability for decoder-only architecture still exists a research gap. We systematically studies whether decoder, an architecture that has shown amazing potential in LLMs, can also take root in learning visual representation through visual supervised training. The key component – causal self-attention we used – is not novel and is inherited from existing LLM architectures, but we propose pivotal techniques, i.e., PS [cls] and soft mask, to effectively adapt them to visual tasks. We hope that this work will inspire more exploration towards a generic pure decoder architecture that fully unifies vision and text.
@@ Appendix A Network Configuration
We provide a block-level comparison between iLLaMA and ViT model in Figure 5. VisionLLaMA uses SwiGLU, and AS2D RoPE to build LLaMA-style architecture. Differently, we further uses RMSNorm, modified causal self-attention and 1D RoPE from LLaMA to replace layer normalization, bi-directional self-attention, and proposes two pivotal strategies, i.e., PS [cls] and soft mask to help the optimization of our iLLaMA. We also keep the learnable positional embedding as ViT does.
@@ B.2 ImageNet (Pre-)training Recipe
We pre-trained the iLLaMA-B/L models on ImageNet-21K for 90 epochs, adhering to the practices in liu2022convnet . We set the cutoff epochs to 30, indicating that the iLLaMA models’ self-attention fully transitions to causal self-attention after 30 epochs.
@@ B.4 Quantization-aware Training Recipe
We provide our quantization-aware training recipe for iLLaMA in Table 15. Basically we follow the Q-ViT method proposed in li2022q , with only weights and activations in each basic block’s causal self-attention and FFN module are quantized to 8 bit width.
@@ Appendix C Designing iLLaMA: detailed results
Table 17: ImageNet-1K classification accuracy via gradually replacing components in ViT-T/16 and ViT-B/16 with counterparts in LLaMA, better or worse than the ViT baseline results with our basic training recipe. Components from or modified from LLaMA are highlighted. P.E.: positional embedding, Bd.: bi-directional self-attention, Cs.: causal self-attention.
@@
Appendix D Rank Analysis of causal Self-attention